OpenStudy (anonymous):

what is the ratio of x to y 3/5=x/x+2 y

OpenStudy (anonymous):

@Baby_Bear69

OpenStudy (anonymous):

what is the ratio of x to y 3/5=x/x+2 y

OpenStudy (baby_bear69):

3/5=x/(x+2y) multiply both side by x+2y (cross multiply) x=3/5 x (x+2y) x=3x/5 + 6y/5 x=(3x+6y)/5 multiply both side by 5 5x=3x+6y subtract both side by 3x 2x=6y simplify both side by 2 x=3y divide both side by y x/y=3

OpenStudy (anonymous):

thank you

OpenStudy (baby_bear69):

\(\large\color{green}{You're~Welcome!!}\)